### v2

# Bargaining via Weber’s law

Abstract
We solve the two-player bargaining problem using Weber’s law in psychophysics, which is applied to the perception of utility changes. By applying this law, one of the players (or both of them) defines lower and upper utility thresholds, such that once the lower threshold is established, bargaining continues in the inter-threshold domain where the solutions are acceptable to both parties. This provides a sequential solution to the bargaining problem, and it can be implemented iteratively reaching well-defined outcomes. The solution is Pareto-optimal, symmetric, and is invariant to affine-transformations of utilities. For susceptible players, iterations are unnecessary, so they converge in one stage toward the (axiomatic) Nash solution of the bargaining problem. This situation also accounts for the asymmetric Nash solution, where the weights of each player are expressed via their Weber constants. Thus the Nash solution is reached without external arbiters and without requiring the independence of irrelevant alternatives. For non-susceptible players our approach leads to different results.
Keywords: Bargaining, Nash Bargaining, Weber’s law, Just noticeable difference
JEL Classification: C71, C78
1 Introduction
Axiomatic bargaining formalizes real-world processes of negotiations, resource allocation, fairness, equity, etc. [1, 2, 3]. Besides social sciences bargaining theories find applications in thermodynamics [4, 5]. The first axiomatic solution of the two-player bargaining problem was deduced by Nash via 4 axioms [1, 2, 3]: (i) Invariance to affine utility transformations. (ii) Symmetry between the players. (iii) Pareto-optimality. (iv) Independence of irrelevant alternatives. Several other solutions have been proposed, e.g. Kalai-Smorodinsky and Kalai egalitarian [3]. They generalize Nash axioms in one way or another.
The drawback of axiomatic bargaining is that it does not provide a constructive framework for realistic negotiations. Zeuthen, and later Harsanyi, studied an iterative scheme of mutual compromises, leading to the Nash solution [1, 6, 7]. Other iterative procedures were presented recently [8, 9]. The first of them converges to the Nash solution, but (like Zeuthen’s theory) external arbiters are required to force players to act accordingly.
We develop a bargaining theory based on Weber’s law (WL) [10, 11]: only those absolute differences $\Delta S=S^{\prime}-S$ between two stimuli $S^{\prime}$ and $S$ are discriminated for which
$|\Delta S|\geq kS,\quad 1>k>0.$ (1)
where $S$ is the intensity of the baseline stimulus, and where $k>0$ is a constant that does not depend on $\Delta S$ and $S$ but depends on the nature of the stimulus and the perceiving individual. Baseline means that the stimulus $S$ came first, and is considered certain, i.e. $S^{\prime}$ is to be discriminated from $S$, and on the background of $S$. The minimal absolute difference $|\Delta S|=kS$ defines the concept of the just-noticeable-difference. Limitations of WL are discussed in [10, 12]. Eq. (1) does not depend on the units chosen to measure $S$. This is crucial for applications of (1) to bargaining.
WL relates to the Weber-Fechner law, which involves additional assumptions, but can be easier to verify in practice [13]. WL has traditionally been applied to the perception of weight, loudness, brightness, and time [10, 11, time_weber]. More recently, the law was observed and validated on cognitive systems, e.g. the perception of temporal durations [14], or representation of numbers and numerosity in humans and animals [15, 16].
Luce applied WL to utility discrimination, which relates to our set-up, and developed utility representation of semi-orders [17, 18]. Argenziano and Gilboa related WL to the Cobb-Douglas utility [19], and applied the law to the concept of interpersonal utility comparison [20]. WL was also employed in price perception [13] and consumer practices (product design, packaging, size of advertisements, frequency of presentation) [21]; see [22] for a critical review.
Using WL, and postulating that the choices are made on the Pareto line, we can define (say for the first player, $\mathcal{P}_{1}$) two utility thresholds, such that when the lower threshold is employed as a baseline stimulus, the inter-threshold interval contains solutions that are acceptable for both players. An alternative formulation of the lower threshold is that $\mathcal{P}_{1}$  has the maximal utility, which can still be increased in a way that is both noticeable to $\mathcal{P}_{1}$  and indifferent to the second player $\mathcal{P}_{2}$. If $\mathcal{P}_{2}$  makes a similar proposal, the resulting iterative solution underlines sequential bargaining: accept modest proposals first, then continue iteratively until the full resource sharing is achieved.
We show that for susceptible players (in Weber’s sense) no iterations are needed, because they effectively converge in one step to the Nash axiomatic solution. When the players hold WL (1) with the same (small) $k$, but have different utilities, we recover the usual Nash solution (maximization of the utility product). For different (but still small) coefficients $k$, we recover for susceptible players the asymmetric Nash solution with the weights given by $k$’s. Hence the Nash solution is reached without postulating the independence of irrelevant alternatives, and without external arbiters. Both conditions were criticized in literature [3, 23, 24].
2 Bargaining set-up
We study two-player bargaining and make the following standard assumptions [2]. $X$ is the amount of a resource, which $\mathcal{P}_{1}$  and $\mathcal{P}_{2}$  bargain to split. $x_{1}$ is the resource $\mathcal{P}_{1}$  gets, and $x_{2}=X-x_{1}$. The utility functions of each player are
$U_{1}(x_{1}),\qquad U_{2}(x_{2}).$ (2)
$U_{1}(x)$ and $U_{2}(x)$ are strictly increasing, bounded from below, smooth, concave, and can be subject to an affine transformation without changing its content; e.g. $U_{1}(x)\to a_{1}U_{1}(x)+b_{1}$, where $a_{1}>0$. Hence we take $U_{1}(x)\geq 0$ and $U_{2}(x)\geq 0$. If no agreement is reached in the bargaining process, the players receive (defection) utilities $d_{1}$ and $d_{2}$. We demand $U_{1}(x_{1})\geq d_{1}$ and $U_{2}(X-x_{1})\geq d_{2}$. The Pareto line is defined via a strictly decreasing, concave, smooth, function $h(\cdot)$:
$u_{2}=h(u_{1}),\qquad h^{\prime}(y)<0,\,h^{\prime\prime}(y)\leq 0,$ (3)
where $h(u_{1})\equiv U_{2}(X-U_{1}^{-1}(u_{1}))$.
As a proposer, each player acts as follows. Our first axiom is that the proposal is made on the Pareto line (3). The second axiom is that $\mathcal{P}_{1}$  and $\mathcal{P}_{2}$  hold (1) for relevant utility changes with constants $k_{1}$ and $k_{2}$, respectively. This means that the utility $u_{1}$ of $\mathcal{P}_{1}$  is maximized in two steps. First $u_{1}-d_{1}$ is taken as the baseline in (1), so that it is still possible to increase it by $w_{1}$. This increase is relevant for $\mathcal{P}_{1}$  (i.e. $\frac{w_{1}}{u_{1}-d_{1}}\geq k_{1}$ according to (1)), and is acceptable for $\mathcal{P}_{2}$(i.e. $\frac{h(u_{1})-h(u_{1}+w_{1})}{h(u_{1})-d_{2}}\leq k_{2}$). The resulting maximal utility $u_{1}^{*}$ is the penultimate noticeable step before hitting the acceptance level of the opponent:
$u_{1}^{*}=\underset{u_{1},\,w_{1}}{{\rm max}}\left[u_{1}\left|~{}~{}\frac{h(u_{1})-h(u_{1}+w_{1})}{h(u_{1})-d_{2}}\right.\leq k_{2};\right.$
$\left.\frac{w_{1}}{u_{1}-d_{1}}\geq k_{1}\right].$ (4)
Theorem 1 below provides an alternative interpretation of (2). $\mathcal{P}_{2}$  makes a similar proposal:
$u_{2}^{*}=\underset{u_{2},\,w_{2}}{{\rm max}}\left[u_{2}\left|~{}~{}\frac{g(u_{2})-g(u_{2}+w_{2})}{g(u_{2})-d_{1}}\right.\leq k_{1};\right.$
$\left.\frac{w_{2}}{u_{2}-d_{2}}\geq k_{2}\right],$ (5)
where $g(.)$ is the inverse of $h(.)$. Eqs. (3) imply:
$y=h(g(y)),\qquad g^{\prime}(y)<0,\,g^{\prime\prime}(y)\leq 0.$ (6)
Figure 1: A representation of the bargaining process with different coefficients $k_{1}$ and $k_{2}$ in (1), Pareto’s line given by $u_{2}=h(u_{1})=\sqrt{X-u_{1}^{2}}$, and the defection point $(0,0)$. Here $u_{1}$ and $u_{2}$ are (resp.) the utilities of $\mathcal{P}_{1}$  and $\mathcal{P}_{2}$, and $X$ is the resource to be shared; cf. (3, 11, 12). The vertical [horizontal] black dashed line shows the result (2) [(2)]. Blue dotted lines are solutions of the second bargaining step; see section 4.
3 One-stage bargaining
Remark 1. Note that (2, 2) are invariant against an affine transformation of utilities. To see this, write (2) as
$u^{*}_{1}=\underset{(u_{1},u_{2}),\,(u_{1}^{\prime},u_{2}^{\prime})}{{\rm max}}\Big{[}u_{1}\Big{|}~{}u_{1}^{\prime}>u_{1},\,u_{2}^{\prime}<u_{2},$
$\frac{u_{2}-u^{\prime}_{2}}{u_{2}-d_{2}}\leq k_{2};~{}~{}\frac{u_{1}^{\prime}-u_{1}}{u_{1}-d_{1}}\geq k_{1}\Big{]},$ (7)
where the maximization goes over the 2 points $(u_{1},u_{2})$ and $(u_{1}^{\prime},u_{2}^{\prime})$ located on the Pareto line and hold $u_{1}^{\prime}>u_{1}$, $u_{2}^{\prime}<u_{2}$.
In (2, 2) we assume $d_{1}=d_{2}=0$ for simplicity. The feasible domain of utilities is $\{u_{1}\geq 0,u_{2}\geq 0\}$ limited by a convex curve; see Fig. 1.
Theorem 1. The problem (3, 2, 1), with $d_{1}=d_{2}=0$ for simplicity, is solved as
$\frac{h(u_{1}^{*}+w_{1}^{*})}{h(u_{1}^{*})}=1-k_{2},\qquad\frac{w_{1}^{*}}{k_{1}}=u_{1}^{*},$ (8)
where $w_{1}^{*}$ is the optimizing value in (2). Eqs. (8) saturate the constraints in (2). The meaning of (8) is that $u_{1}^{*}$ and $u_{1}^{*}+w_{1}^{*}$ can be considered as upper and lower utility thresholds. If $u_{1}^{*}$ is stated by $\mathcal{P}_{1}$  and accepted as the baseline, the utility interval $(u_{1}^{*},u_{1}^{*}+w_{1}^{*})$ is acceptable for both players. Increasing the utility beyond $u_{1}^{*}+w_{1}^{*}$ is noticeable for $\mathcal{P}_{1}$  and non-acceptable for $\mathcal{P}_{2}$.
Proof. Write conditions in (2) as
$\frac{h(u_{1}+w_{1})}{h(u_{1})}\geq 1-k_{2},\qquad\frac{w_{1}}{u_{1}}\geq k_{1}.$ (9)
Let $F_{w_{1}}(y)$ be the inverse function of $y(u_{1})=\frac{h(u_{1}+w_{1})}{h(u_{1})}$. Eqs. (3) imply $y^{\prime}(u_{1})<0$ hence $\frac{{\rm d}F_{w_{1}}}{{\rm d}y}\equiv F^{\prime}_{w_{1}}(y)<0$. Then (9) reads
$u_{1}^{*}=\underset{w_{1}}{{\rm max}}[\,{\rm min}[F_{w_{1}}(1-k_{2}),\,w_{1}/k_{1}]\,].$ (10)
Using (1), this equation implies $F_{w_{1}^{*}}(1-k_{2})=w_{1}^{*}/k_{1}$, where $w_{1}^{*}$ is the optimizer in (10). This leads to (8).
Example. Theorem 1 solves (2) explicitly for certain functions $h(x)$. Take for example in (2):
$U_{1}(x_{1})=x_{1}^{p},~{}~{}U_{2}(x_{2})=x_{2}^{q}=(X-x_{1})^{q},$ (11)
$u_{2}=h(u_{1})=\left(X-u_{1}^{1/p}\right)^{q},~{}~{}0<p,q<1.$ (12)
Ref. [25] presents empirical support for utility functions (11) and estimates parameters $p$ and $q$. Eqs. (1, 2, 8) have the following solutions:
$(u_{1}^{*})^{1/p}=\epsilon_{1}X,\quad(u_{2}^{*})^{1/q}=\epsilon_{2}X,$ (13)
$w_{1}^{*}=k_{1}\epsilon_{1}^{p}X^{p},\qquad w_{2}^{*}=k_{2}\epsilon_{2}^{q}X^{q},$ (14)
$\epsilon_{1}=\frac{1-(1-k_{2})^{1/q}}{(1+k_{1})^{1/p}-(1-k_{2})^{1/q}},$ (15)
$\epsilon_{2}=\frac{1-(1-k_{1})^{1/p}}{(1+k_{2})^{1/q}-(1-k_{1})^{1/p}},\qquad 0<\epsilon_{1},\epsilon_{2}<1.$ (16)
According to (16, 13), the amount of resource obtained by each player depends on the utilities and $k$-coefficients of both players. Note that
$x_{1}^{*}+x_{2}^{*}\leq X,$ (17)
where the equality in (17) is reached only for $p=q=1$. Hence apart from this case, the resource is not exhausted. This happens approximately whenever $k_{1}$ and $k_{2}$ in (16) are sufficiently small:
$\epsilon_{1}\simeq\frac{(k_{2}/q)}{(k_{2}/q)+(k_{1}/p)},\qquad\epsilon_{2}\simeq\frac{(k_{1}/p)}{(k_{2}/q)+(k_{1}/p)}.$ (18)
Remark 2. A particular case of (8) coincides with the Nash solution. Take $k_{1}$ and $k_{2}$ sufficiently small, but their ratio $k_{1}/k_{2}$ is finite; cf. (18). We Taylor expand in (8): $h(u_{1}^{*}+w_{1}^{*})\simeq h(u_{1}^{*})+w_{1}^{*}h^{\prime}(u_{1}^{*})$. Then (8) and its analog for $\mathcal{P}_{2}$  produce
$\frac{h^{\prime}(u_{1}^{*})u_{1}^{*}}{h(u_{1}^{*})}=-\frac{k_{2}}{k_{1}},\qquad\frac{g^{\prime}(u_{2}^{*})u_{2}^{*}}{g(u_{2}^{*})}=-\frac{k_{1}}{k_{2}}.$ (19)
The first implication of (19) is that $u_{2}^{*}=h(u_{1}^{*})$, as seen from using the derivative of the inverse function in one of equations in (19). The second implication is that $u_{2}^{*}=h(u_{1}^{*})$ coincides with the Nash solution:
$u_{1}^{*}={\rm argmax}_{u}\left[u^{\frac{k_{2}}{k_{1}+k_{2}}}h(u)^{\frac{k_{1}}{k_{1}+k_{2}}}\right],$ (20)
where (20) is proved by differentiation, and where $u_{1}^{\frac{k_{2}}{k_{1}+k_{2}}}h(u_{1})^{\frac{k_{1}}{k_{1}+k_{2}}}=u_{1}^{\frac{k_{2}}{k_{1}+k_{2}}}u_{2}^{\frac{k_{1}}{k_{1}+k_{2}}}$ is precisely the weighted product of utilities that appears in the (asymmetric) Nash solution [2]. Importantly, the axiomatic Nash solution acquires psychophysical interpretation. Note that a more susceptible player ($\mathcal{P}_{1}$  if $k_{1}<k_{2}$) has a larger weight in (20). This does not mean that a more susceptible player necessarily gets more resources; see (18).
Theorem 2. For (1) and a strictly decreasing and strictly concave $h(\cdot)$, the following inequalities are true
$h(u_{1}^{*}+w_{1}^{*})<u_{2}^{*}<h(u_{1}^{*}),$ (21)
$g(u_{2}^{*}+w_{2}^{*})<u_{1}^{*}<g(u_{2}^{*}).$ (22)
Remark 3. Eqs. (21, 22) mean that none of the players should make the last step (e.g., $u_{1}^{*}\to u_{1}^{*}+w_{1}^{*}$ should not be demanded) since $\mathcal{P}_{1}$  will demand more than $\mathcal{P}_{2}$  envisaged for $\mathcal{P}_{1}$: $u_{1}^{*}+w_{1}^{*}>g(u_{2}^{*})$; i.e., cooperation demands not making the last step.
Proof. We prove only (21). Rewrite (8) as
$\frac{h(\,[1+k_{1}]u_{1}^{*})}{h(u_{1}^{*})}=1-k_{2},$ (23)
$\frac{g(\,[1+k_{2}]u_{2}^{*})}{g(u_{2}^{*})}=1-k_{1}.$ (24)
Eqs. (3, 6) imply that $\frac{h((1+k_{1})u_{1})}{h(u_{1})}$ and $\frac{g((1+k_{2})u_{2})}{g(u_{2})}$ are decreasing functions of $u_{1}$ and $u_{2}$, respectively.
To prove (21), we assume $u_{2}^{*}\geq h(u_{1}^{*})$. Using this with the second equality in (24), we get
$1-k_{1}=\frac{g(\,[1+k_{2}]u_{2}^{*})}{g(u_{2}^{*})}\leq\frac{g(\,[1+k_{2}]\,h(u_{1}^{*})\,)}{u_{1}^{*}},$ (25)
$(1+k_{1})u_{1}^{*}\leq g(\,[1+k_{2}]\,h(u_{1}^{*})\,).$ (26)
With $h(\cdot)$ decreasing we get from (26)
$h([1-k_{1}]u_{1}^{*})\geq(1+k_{2})h(u_{1}^{*}),$ (27)
$h([1+k_{1}]u_{1}^{*})=(1-k_{2})h(u_{1}^{*}).$ (28)
Multiplying both sides of (27) by $\frac{1}{2}$, adding them, and using the strong concavity of $h(\cdot)$, we get a contradiction:
$h(u_{1}^{*})>\frac{1}{2}h([1-k_{1}]u_{1}^{*})+\frac{1}{2}h([1+k_{1}]u_{1}^{*})\geq h(u_{1}^{*}).$ (29)
Contrary to (21) we assume:
$u_{2}^{*}\leq h(\,[1+k_{1}]\,u_{1}^{*}).$ (30)
Plugging this into (24):
$\frac{g\left(\,\,[1+k_{2}]h([1+k_{1}]u_{1}^{*})\,\,\right)}{[1+k_{1}]u_{1}^{*}}\leq(1-k_{1}),$ (31)
$(1+k_{2})h(\,[1+k_{1}]u_{1}^{*}\,)\geq h(\,[1-k_{1}^{2}]\,u_{1}^{*}),$ (32)
$(1+k_{2})h([1+k_{1}]\,u_{1}^{*})=(1-k_{2}^{2})\,h(u_{1}^{*}),$ (33)
where (33) follows from multiplying (23) by $(1+k_{2})$. Eqs. (32, 33) produce
$(1-k_{2}^{2})\,h(u_{1}^{*})\geq h(\,[1-k_{1}^{2}]\,u_{1}^{*}).$ (34)
This contradicts to the monotonous decay of $h(.)$ unless $k_{1}=k_{2}=0$, which contradicts (1).
4 Iterative bargaining
Eq. (17) and Theorem 2 show that within the one-stage bargaining, $\mathcal{P}_{1}$  and $\mathcal{P}_{2}$  insist on $u_{1}^{*}$ and $u_{2}^{*}>h(u_{1}^{*})$, i.e. the resource is not exhausted. Since utilities are not additive, we translate $u_{1}^{*}$ and $u_{2}^{*}$ to corresponding resources, subtract them from the full resource, and consider the remaining resource as the subject of bargaining in the second round. Rounds converge as the example (1116) illustrates. Given (13), the new resource for the second stage of bargaining will be $(1-\epsilon_{1}-\epsilon_{2})X$, and the resources acquired by (resp.) $\mathcal{P}_{1}$  and $\mathcal{P}_{2}$  in the second stage will be $\epsilon_{1}(1-\epsilon_{1}-\epsilon_{2})X$ and $\epsilon_{2}(1-\epsilon_{1}-\epsilon_{2})X$. Increasing the iteration steps to infinity, we find that the resources obtained (resp.) by $\mathcal{P}_{1}$  and $\mathcal{P}_{2}$  read:
$\epsilon_{1}X\sum_{m=0}^{\infty}(1-\epsilon_{1}-\epsilon_{2})^{m}=\frac{\epsilon_{1}X}{\epsilon_{1}+\epsilon_{2}},\qquad\frac{\epsilon_{2}X}{\epsilon_{1}+\epsilon_{2}}.$ (35)
Hence the resource is now fully shared, the corresponding weights being $\frac{\epsilon_{1}}{\epsilon_{1}+\epsilon_{2}}$ and $\frac{\epsilon_{2}}{\epsilon_{1}+\epsilon_{2}}$. For the symmetric situation, $k_{1}=k_{2}$, we get that the resource is shared equally due to $\epsilon_{1}=\epsilon_{2}$; see (16). Hence the symmetry axiom holds.
Eqs. (35) are different from the Nash axiomatic solution when $k_{1}=k_{2}$ is not small, but $p\not=q$. The Nash solution is deduced from ${\rm max}_{x_{1}}[x_{1}^{p}(X-x_{1})^{q}]$ and leads to $x_{1}=pX/(p+q)$, which coincides with (35) only for $p=q$.
There are two possible scenarios of bargaining: (i) Players start with exaggerated demands and lower them during bargaining. (ii) Players act sequentially, first making offers that are acceptable for both, and then bargaining over the rest. Thus our results point towards the second scenario; see also Theorem 1.
Acknowledgements
This work was supported by the Higher Education and Science Committee of Armenia, grant No. 21AG-1C038. We acknowledge useful discussions with Lida Aleksanyan and Sanasar Babajanyan.